OpenStudy (anonymous):

write the infinite geometric series in sigma notation. then find the sum, if it exists. show work. the sum of an infinite geometric series is 125, and the value of r is 0.4. find the first three terms of the series.

OpenStudy (amistre64):

the other person gave you a suitable answer to this .....

OpenStudy (amistre64):

either participate in the solution process, or go someplace else please. Reposting is not considered a viable option and is considered "just wanting answers".

OpenStudy (anonymous):

First of all I don't just want the answers but I do want more than one person to help me with the problem. if you refer back to it she wasn't 100% sure if she was right that's why she said she "thinks" she was. so for your information I just wanted more people to comment on it so that I can understand it. thank you very much. @amistre64

OpenStudy (amistre64):

then what questions do you have about her process? it was correct for a start

OpenStudy (amistre64):

r < 1 , so itll converge a/(1-r) is the sum of an infinite series that converges therefore a = 125(1-r)

OpenStudy (anonymous):

well thank you, I just wanted to make sure the process was correct

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

only thing missing i think is the summation for it .....

OpenStudy (amistre64):

since we are looking at an infinite sum, its going to have to involve a limit statment

OpenStudy (amistre64):

\[S_\infty=\lim_{n\to \infty}\frac{a_1(1-r^n)}{1-r}\] you do know how to find the first 3 terms: a1, a2, and a3 right?

OpenStudy (amistre64):

pfft, that a correct equation, but doesnt use summation notation does it \[\lim_{n\to \infty}\sum_{k=0}^{n}a_1~r^{k}\]

OpenStudy (amistre64):

some texts find it ok to put the infinity sign up top of the big E, sigma

OpenStudy (anonymous):

@amistre64 no I am not sure how to go about finding the first 3 terms

OpenStudy (anonymous):

and it needs to be in sigma notation